OpenStudy (anonymous):

A farmer wants to enclose a rectangular field along a river on three sides. If 4,800 feet of fencing is to be used, what dimensions will maximize the enclosed area

OpenStudy (anonymous):

We know that: \[ 2l+2w=4800 \]And we wish to maximize: \[ lw=A \]So, we write one in terms of the other: \[ l+w=2400\implies\\ w=2400-l\implies\\ A=l(2400-l) \]To find the maximum, we can simply find the vertex, but I'll do it a slightly different way here: \[ A=2400l-l^2\implies\\ \frac{d}{dl}A=2400-2l\implies\\ \]A is a maximum when A'=0, so: \[ 0=2400-2l\implies\\ l=1200\\ w=2400-l=2400-1200=1200 \]Therefore the maximum is, when \(l=1200\) and \(w=1200\) Of course, this problem can be simplified to finding the largest square, but that won't always be the case for these kinds of problems.

OpenStudy (anonymous):

one small thing, i think the correct equation is 2w + l = 4800 for perimeter, but your answer is correct 

OpenStudy (anonymous):

err actually i think length should be.. 2400

OpenStudy (anonymous):

Yes, you are correct. Wow, I feel stupid, for not noticing that... but thank you for the correction!

OpenStudy (anonymous):

I'll re-trace this problem quickly, to correct my mistake: \[l+2w=4800\\ l=4800−2w\\ lw=A\] Which means: \[A=w(4800−2w)=4800w−2w^2\implies\\ A′=4800−4w\]Gives a maxima at:\[ 4w=4800\implies\\ w=1200 \]Which means the length is:\[ l=4800−2(1200)=2400 \]Thus: \[ l=2400\\ w=1200\] Jeez, I screwed up twice.. need to get some sleep! G'night!

OpenStudy (anonymous):

Thanks, GN